# Towards Reliable Latent Knowledge Estimation in LLMs: In-Context Learning vs. Prompting Based Factual Knowledge Extraction

### v1

@@ 2.2 A new In Context learning based LKE (IC-LKE)
Our goal is to estimate whether an LLM knows a fact . The challenge is to probe the LLM and evaluate its responses in a way compatible with the design principles set in Section 2.1.
@@
Key idea: Leverage in-context learning. LLMs have shown to exhibit In-Context Learning (ICL) abilities (Brown et al., 2020) that allow them to infer and extrapolate patterns in their inputs. We leverage this ability to communicate information about relation without additional instructions to the LLM (DP3) by providing it with a list of facts based on .
@@
Assume that we want to probe for whether an LLM knows the fact Einstein, birth-year, 1879 . We can use other facts for the birth-year relation such as Feynman, birth-year, 1918 Heisenberg, birth-year, 1901 to construct an input “Feynman 1918 Heisenberg 1901 Einstein”. By providing in-context examples to the model, we communicate the relation between subjects and objects. To correctly extrapolate the pattern, the model needs to retrieve Einstein’s birth-year as the completion of the sequence.
@@
We use to pick facts from and concatenate the tokens corresponding to the subjects and objects, but do not include any other information about (DP2). We use space “ ” as the separator token and discuss this choice in more detail in Section 4.1. We discuss other design choices for IC-LKE construction in Section 3. When further details are not needed, we simply refer to some input as .
Evaluating model outputs. We evaluate the output of model for input based on the probabilities assigns to the tokens of the corresponding object . To allow for objects consisting of multiple tokens and to be independent of the specific tokenization scheme (DP1), we compute the object probability over multiple tokens as follows:
where denotes the number of tokens in and is the conditional probability of predicting the -th token of given the preceding tokens , and .
Multiple-choice testing. To determine whether model knows a fact , we test whether given input , can choose the correct object from among a set of unique alternatives. Specifically, given fact , we derive a test instance called choice , where is a set of plausible but incorrect alternatives. We discuss the choice of in Section 4.
denotes the prediction of for choice . The predicted object has the maximal object probability within .
Evaluation Metric. We evaluate the factual knowledge of model over a dataset of choices using multiple choice accuracy:
@@
The IC-LKE design satisfies the knowledge estimation design principles. The IC-LKE design proposed here satisfies the design principles from Section 2.1, since
DP1: its relative probability comparisons between different answer-options make it applicable to arbitrary types of facts.
DP2: it uses the same, minimal prompt design based on ICL across all relations.
DP3: its only requirement is that the LLM is able to use ICL, no further assumptions about any metalinguistic abilities are made.
3 Exploring the design space of IC-LKE
examples in a prompt
unknown examples
unknown examples
incorrect examples
incorrect examples
By design, IC-LKE avoids many limitations of prior works. However, IC-LKE introduces a few design choices for the input, i.e., in Equation (1). One must decide the right , the number of in-context examples included in . Further, it is unclear how IC-LKE would be impacted when some of the chosen examples are unknown to the model or are incorrect. We study both these factors in detail by varying and introducing unknown or incorrect examples within these examples. These experiments allows us to better understand the number of in-context examples needed and how robust IC-LKE is to several types of noise in these in-context examples. We perform an in-depth empirical analysis on a Nobel Laureate dataset for the relation ‘birth year’ (details in A.1). The dataset consists of facts formatted as .
@@ 3 Exploring the design space of IC-LKE
More knowledgeable models need fewer in-context examples, but a small number suffices for most models. In Figure 1, we report knowledge estimation accuracy (Eq. (5)) for different LLMs evaluated on 900 test samples, with varying numbers of in-context examples () by randomly sampling from the training set using five random seeds. With an increasing number of in-context examples, the mean accuracy increases while the standard deviation decreases in different LLMs, i.e., the models gradually converge to a stable performance. Using dashed vertical lines, we report the minimum number of examples required by different LLMs to achieve of the accuracy at in-context examples. Interestingly, LLMs with higher estimation accuracy tend to require fewer in-context examples compared to those with lower accuracy. A potential explanation for this behavior is that in order to infer the relation , models need to comprehend the examples presented in the prompt. Therefore, less knowledgeable models need to see more examples in order to infer . To further investigate which individual facts may be known or unknown to a model, we look at the generation probability of in-context objects in correct subject ()-object () pairs using the Mistral-7B model, as shown in Figure 2a. Similar results for additional models are presented in Appendix E. Note that here we are only looking at probabilities of the object () for in-context examples given previous pairs in the input to understand which of these samples are known by the LLM. The Mistral-7B model demonstrates a gradual increase in probability for generating correct objects as we go from left to right on the x-axis (note that for a point on the x-axis, points before it are in context, thus points on the right have more context to leverage) in Figure 2a, stabilizing at a mean probability of approximately 85%. We also see that some objects at later positions have a lower generation probability. This suggests that the LLM may be less confident about its knowledge of the facts corresponding to them. We can leverage the token generation probability as a signal of LLM’s confidence when evaluating LKEs (see Appendix D).
@@
Models are robust to unknown examples. Next, we investigate the robustness of estimates to occurrence of unknown examples. We insert unknown examples in two distinct ways: one where we randomly distribute the occurrence of unknown examples throughout , and another more extreme scenario where we replace a continuous block of examples with unknown ones. We chose out of the examples and replaced them with unknown examples created using fictitious names and birth years 111generated via https://en.namefake.com/api. Our findings are shown in Figures 2b and 2c for random and continuous replacement respectively. Unknown examples are marked by red dots, examples immediately following unknown ones in cyan dots and the rest in blue dots. The unknown examples show generation probabilities close to zero, confirming the LLM’s tendency to assign low probabilities to unknown data. However, interestingly, unknown examples minimally impact surrounding data in both settings.
@@
Models are vulnerable to incorrect examples. We investigate the impact of including incorrect examples in . Similar to the setup for unknown examples, we also insert 40 (out of 200) incorrect examples randomly (Figure 2d) and simultaneously (Figure 2e). In our experiments, these incorrect examples are created by altering the birth years of known Nobel laureates and are marked by red dots in the plots. In contrast to inserting unknown examples, the LLM significantly struggles with incorrect examples. Injection of such examples detrimentally affects the LLM’s performance in both settings. We highlight one randomly marked yellow star example in Figure 2a, Figure 2b, and Figure 2d to show how the presence of incorrect samples brings down the probability of surrounding points.
Summary: LLMs can identify the relation pattern of subject-object pairs even with a small set of in-context examples in the prompt. LLMs are relatively robust to unknown examples, but their ability to recollect factual knowledge is vulnerable to incorrect examples, particularly when they appear in a continuous sequence. Our findings allude to the effectiveness of designing an IC-LKE, where we carefully place correct examples from a training dataset and proceed to estimate the latent knowledge of the LLM on examples from the test set. Furthermore, the findings also motivate us to design a more efficient in-context learning based LKE, called EIC-LKE, that can process multiple test examples simultaneously in a single prompt where training examples are placed preceding each test example, see more details in the Appendix F.
4 Experiments and Results
We present the empirical findings of IC-LKE (as well as the efficient version, EIC-LKE) on the knowledge-estimation task on 49 open-source (pre-trained and fine-tuned) LLMs across different LLM families and sizes. We enlist models and their simplified names used in this paper in Appendix 6, Table 6, and provide a leader-board of models based on IC-LKE in Table LABEL:table:model_order.
@@ 4 Experiments and Results
Dataset: We evaluate the knowledge of models on a large set of facts from the T-REx dataset222https://huggingface.co/datasets/relbert/t_rex Elsahar et al. (2018). We selected relations from T-REx with at least 500 samples and linked to a minimum of 100 unique objects. This filtering leads to 50 distinct relations spanning categories like birth dates, directorial roles, parental relationships, and educational lineage. The resulting T-REx Multiple Choice (T-REx-MC) dataset comprises 5,000 training and 20,000 test facts. Appendix A contains detailed information on the dataset and relations.
@@
4.1 IC-LKE vs. prompt-based approaches
@@
We compare the performance of IC-LKE and EIC-LKE with the existing prompt-based approaches (Jiang et al., 2020) and report two key takeaways.
IC-LKE outperforms prompt-based approaches. We randomly sample three human-generated prompts (HGP) and machine-mined prompts (MMP) from (Jiang et al., 2020) for 12 common relations between T-REx-MC and (Jiang et al., 2020). The HGPs and MMPs for all relations are in Appendix G. In Figure 3, IC-LKE and EIC-LKE outperform HGP and MMP in terms of higher mean accuracy across different models and 12 relations. Also, IC-LKE and EIC-LKE have lower standard deviation than HGP and MMP, indicating a higher consistency of IC-LKE and EIC-LKE on knowledge estimation tasks. In Appendix H.2, we report relation specific results, where IC-LKE and EIC-LKE estimate higher factual knowledge than the existing works in most relations, thereby demonstrating the superiority of IC-LKE and EIC-LKE over existing methods.
@@
IC-LKE is a flexible and effective knowledge estimator. We adapt IC-LKE by replacing the separator ‘[space]’ with three separators from HGP and MMP each for the relation ‘original broadcaster’ and report estimation accuracy in Figure 4. We can observe that ‘[space]’ token demonstrates an equivalent performance with semantically meaningful prompts via HGP and MMP. Therefore, adding relation specific separators has a limited impact on factual knowledge estimation, as long as the subject-object pairs are correctly presented. Furthermore, finding relation-specifc prompts often require hand-crafted efforts vs. an automatic in-context based approach like ours where (subject, object) pairs are used. Therefore, IC-LKE can potentially extend to any facts from knowledge graphs over any LLM while HGP and MMP requires additional supervision and relation-specific validation.
4.2 Evaluating Diverse Models and Relations
@@
We investigate the performance of 35 pre-trained LLMs and 14 fine-tuned LLMs across 50 relations using the IC-LKE framework. Our analysis is designed to uncover nuanced insights into the knowledge levels and structures within these models. We will examine the results through two primary lenses: (1) the variations in knowledge across different model families, and (2) the influence of model size and fine-tuning within the same model family on their knowledge attributes.
4.2.1 Comparing different LLMs families
@@
Some model families are consistently more knowledgeable than the rest. We sort the model families based on the performance of the model closest to 7B parameters 3337B parameters is a good reference point since all model families except GPT-NEO-X have models within a gap of 1B parameters: Mistral-7B, Gemma-7B, Llama-7B, Falcon-7B, MPT-7B, OPT-6.7B, GPT-J-6B, Pythia-6.9B, and Bloom-7.1B., and the models within each family based on average accuracy across 50 relations. Figure 5 shows that the Mistral, Llama2, Gemma, and Llama families have higher performance on most of the relations than Pythia, Bloom, and OPT, indicating their lower factual knowledge.
@@
Different model families align in their relative factual knowledge. We investigate the correlations between each model pair’s performance over 50 relations to assess the agreement in their knowledge levels of the 50 relations. We compute the average correlations within each model family (e.g. Llama2 7B, 13B, 70B) in Figure 6. Despite differences in architecture and training datasets among model families, there is a significant consensus (correlation > 0.6, see Figure 14) regarding the hierarchy of knowledge across various relations. We also compile the three best and worst-performing relations for each model in Table 9, illustrating the consensus among all models.
4.2.2 Comparing within the same LLM family

### v2

@@ 2.2. A new Zero- Prompt based LKE (ZP-LKE)
Our goal is to estimate whether an LLM knows a fact . The challenge is to probe the LLM and evaluate its responses in a way compatible with the design principles defined in Section 2.1.
@@
The key idea here is to eliminate prompts meant to capture the relation (zero-prompt) and instead rely on examples of similarly related pairs to probe the internal knowledge. LLMs have been shown to exhibit In-Context Learning (ICL) abilities (Brown et al., 2020) that allow them to infer and extrapolate patterns in their inputs. We leverage this ability to communicate information about relation without additional instructions to the LLMs (DP1 and DP2) by providing it with a list of facts based on .
@@
Assume that we want to probe for whether an LLM knows the fact Einstein, birth-year, 1879 . We can use other facts for the birth-year relation such as Feynman, birth-year, 1918 Heisenberg, birth-year, 1901 to construct an input “Feynman 1918 Heisenberg 1901 Einstein”. By providing such zero-prompt in-context examples to the model, we expect to communicate the underlying relation between subjects and objects. To correctly extrapolate the pattern, the model needs to retrieve Einstein’s birth-year as the completion of the sequence.
@@
We use to pick facts from and concatenate the tokens corresponding to the subjects and objects, but do not include any other information about . We use space “ ” as the separator token and discuss this choice in detail in Section 4.1. We discuss other design choices for the construction of ZP-LKE in Section 3. When further details are not needed, we simply refer to some input as .
ZP-LKE design satisfies all our design principles (DPs):
DP1: by construction, zero-prompting eliminates prompt hacking and thus, risks of over-fitting and side-channels.
DP2: it only relies on the in-context learning abilities and not meta-linguistic judgments of an LLM.
DP3: by construction, input is LLM-agnostic and hence, enables cross-LLM latent knowledge comparisons.
2.3. Evaluating model outputs
We evaluate the output of model for input in two ways: (1) Open-ended generation that lets the model generate till tokens (Kalo and Fichtel, [n. d.]; Yu et al., 2024) after which the presence of the ground truth is checked within the response, (2) Multiple-choice test that forces the model to predict from a list of options (Jiang et al., 2020).
(1) Response testing in open-ended generation. Given a fact and a model , we provide the input to the model and let it generate for tokens . We consider the answer to be correct if leading to the prediction .
(2) Multiple-choice testing. In the multiple-choice testing, we extract the answer based on the probabilities assigned to the tokens of the corresponding object . To allow for objects consisting of multiple tokens and to be independent of the specific tokenization scheme or LLM (DP3), we compute the object probability over multiple tokens as follows:
where denotes the number of tokens in and is the conditional probability of predicting the -th token of given the preceding tokens , and . To determine whether model knows a fact , we test whether given an input , can choose the correct object from among a set of unique alternatives. Specifically, given fact , we redefine it as , where is a set of plausible but incorrect alternatives. We discuss the choice of in Section 4.
denotes the prediction of for the fact . The predicted object has the maximal object probability within .
We evaluate the factual knowledge of model over a dataset of test facts using accuracy as a metric for both the response test and multiple-choice test:
@@ 2.3. Evaluating model outputs
3. Exploring the design space of ZP-LKE
Our ZP-LKE design avoids many of the reliability concerns of prior works (Jiang et al., 2020; Kalo and Fichtel, [n. d.]; Yu et al., 2024; Hogan et al., 2021; Cao et al., 2021; Petroni et al., 2019). However, ZP-LKE also introduce a few design choices for the input i.e., in Equation (1). One must decide the right , the number of in-context examples included in . Furthermore, it is unclear how ZP-LKE would be affected if some chosen examples are unknown to the model or are incorrect or appeared in a different order.
We study this by varying and introducing unknown or incorrect examples within these examples. While many prior works investigated the number of in-context examples needed for various tasks (Brown et al., 2020; Agarwal et al., 2024; Chen et al., 2023; Lin and Lee, 2024; Pan, 2023), it is worth re-examining them for ZP-LKE for three reasons: (i) prior works report differing results, with some reporting that increasing the number of examples improves performance (Agarwal et al., 2024), while others argue the opposite (Lin and Lee, 2024), (ii) most don’t carefully distinguish between the two learning modes of in-context learning (as noted in Section 1, ZP-LKE relies on one of the modes), and (iii) only a few (Min et al., 2022; Chen et al., 2023) studied the influence of incorrect examples and none studied the impact of unknown examples.
Our experiments help us understand the number of in-context examples needed, as well as how the in-context example’s generation probability changes with different types of noise. We perform an in-depth empirical analysis on a Nobel Laureate dataset for the relation ‘birth year’ (details in Appendix A.1). The dataset consists of facts formatted as .
The number of required in-context samples for communicating both the question and answer format varies across LLMs. In Figure 3, we report multiple choice accuracy (Eq. (4)) for different LLMs evaluated on 900 test samples, with varying numbers of in-context examples () that are randomly sampled from a separate training set using 5 random seeds. As the number of in-context examples increases, the mean accuracy rises while the standard deviation decreases across different LLMs, indicating that the models gradually converge to stable performance.
The dashed vertical lines show the minimum number of examples required by different LLMs to achieve 95% of the accuracy reached with 50 in-context examples. Interestingly, LLMs with higher estimation accuracy require fewer in-context examples than those with lower accuracy to effectively interpret the underlying question. This maybe attributed to the amount of internal knowledge contained in the LLMs. To enable ZP-LKE across all the LLMs, we set for the following experiments.
@@ 3. Exploring the design space of ZP-LKE
We delve deeper to further investigate which individual facts may be known or unknown to a model. We examine the generation probability of in-context objects in 200 correct subject ()-object () pairs using the Mistral-7B model. We can see in Figure 4(a)that Mistral-7B model shows a gradual increase in the probability of generating correct objects from left to right on the x-axis (where points on the right have more context to leverage) stabilizing at a mean probability of approximately 0.85. Some objects at later positions, however, have a lower generation probability, suggesting that the LLM may be less confident about its knowledge of the facts corresponding to those pairs. Thus, we can leverage the generation probability as a signal of the LLM’s confidence when evaluating LKEs (see Appendix D). Similar results for additional models are presented in Appendix E.
@@
Models are robust to unknown examples. Next, we investigate the robustness of estimates to the occurrence of unknown examples. We insert unknown examples in two distinct ways: randomly distributed throughout and in a more extreme scenario, where a continuous block of examples is replaced with unknown ones. We selected 40 out of the 200 examples and replaced them with unknown examples created using fictitious names and birth years 222generated via https://en.namefake.com/api. Our findings are shown in Figures 4(b) and 4(c) for distributed and continuous replacement respectively. Unknown examples are marked by red dots, examples immediately following unknown ones in cyan dots and the rest in blue dots. The unknown examples show generation probabilities close to zero, confirming the LLM’s tendency to assign low probabilities to unknown data. However, interestingly, unknown examples minimally impact the generation probability of the surrounding data in both settings.
@@
Models are vulnerable to incorrect examples. Similar to the setup for unknown examples, we also insert 40 (out of 200) incorrect examples randomly (Figure 4(d)) and simultaneously (Figure 4(e)). In our experiments, these incorrect examples are created by altering the birth years of known Nobel laureates and are marked by red dots in the plots. In contrast to inserting unknown examples, the LLM significantly struggles with the injection of incorrect examples. It detrimentally affects the LLM’s performance in both settings thus revealing the vulnerability of the models towards incorrect examples. We highlight one randomly marked yellow star example in Figure 4(a), Figure 4(b), and Figure 4(d) to show how the presence of incorrect samples significantly brings down the probability of the neighboring points.
Summary: The key takeaways while exploring the design space of ZP-LKE are - (a) Different LLMs take varying numbers of in-context samples to comprehend both the question and format of the answer alongside, with being an optimal number for our setup. (b) The models are robust to unknown examples but vulnerable to incorrect examples. To the best of our knowledge, we are the first to distinguish between examples unknown to the model and incorrect examples known to the model and study their impact on in-context learning. As ZP-LKE relies on many examples, this distinction and understanding is important in practice. Also, while (Agarwal et al., 2024) found that the order of examples has varying effects in different domains, we identify the distribution of unknown and incorrect examples as a crucial underlying factor.
4. Experiments and Results
As ZP-LKE inputs are model-agnostic and easy to adapt for a large variety of relations, it can be used to very effectively to conduct cross-LLM latent knowledge comparisons. We leverage ZP-LKE to estimate latent knowledge across 49 open-source (pre-trained and fine-tuned) LLMs, spanning different LLM families (Llama (2), Mistral, Mixtral, Gemma, Falcon, Pythia, Bloom, and OPT) and sizes (from 70M to 822B). To the best of our knowledge, we are the first to evaluate a knowledge estimation framework across a large number of models. We list the models and their simplified names used in this paper in Appendix G, Table 7, and provide a leaderboard of models based on ZP-LKE in Appendix G, Table LABEL:table:model_order. We hope our results and the framework can help future LLM developers reliably and efficiently estimate the latent knowledge of their models.
@@ 4. Experiments and Results
Dataset: We evaluate the knowledge of models on a large set of facts from the T-REx dataset (Elsahar et al., 2018). We selected relations from T-REx with at least 500 samples that are linked to a minimum of 100 unique objects. We create a list of multiple choices for each sample and ensure that instances with multiple correct objects do not have any of their correct answers in their multiple-choice list. This filtering leads to 50 distinct relations spanning categories like birth dates, directorial roles, parental relationships, and educational lineage. The resulting T-REx Multiple Choice (T-REx-MC) dataset comprises 5,000 training and 20,000 test facts. Appendix A contains detailed information on the dataset and relations.
@@
4.1. ZP-LKE vs. prompt-based approaches
@@
We compare the performance of ZP-LKE with the existing prompt-based approaches (Jiang et al., 2020) using both the response accuracy and the multiple-choice accuracy defined in Section 2.2.
ZP-LKE outperforms prompt-based approaches. We randomly sample three human-generated prompts (HGP) and machine-mined prompts (MMP) from (Jiang et al., 2020) for 12 common relations between T-REx-MC and (Jiang et al., 2020). We show that ZP-LKE outperforms HGP and MMP in terms of the accuracy measures by a large margin, across different models and 12 relations in Figure 5; the detailed accuracy for each relation can be found in the Appendix G.2, Figure 14 and 15.
Figure 5(a) shows that ZP-LKE improves the fraction of facts accurately extracted from four open-source models by an average of 35% for HGPs (from 0.45 to 0.61) and 90% for MMPs (from 0.32 to 0.61). In the case of multiple-choice accuracy, having controlled the influence of the answer format, we observe that all the knowledge estimation methods improve in their performance. Alongside, ZP-LKE still outperforms existing approaches by an average of 9.41% for HGPs (from 0.71 to 0.78) and 57% for MMPs (from 0.50 to 0.78). The multiple-choice accuracy metric disentangles the answer format from the question leading to better factual knowledge estimation across the board. Hence, we primarily report the multiple-choice accuracy metric for the experiments in the rest of the paper.
@@
ZP-LKE performs better than FS-LKE with the same number of examples. We adapt ZP-LKE by replacing the separator token “ ” between subjects and objects with three prompts each from HGPs and MMPs for the relation ‘original broadcaster’ and report the multiple choice accuracy in Figure 6. We intend to understand whether the additional prompt tokens can help in communicating the question better. The ZP-LKE with “ ” token performs equally or even better for some models compared to the semantically meaningful prompts from HGP and MMP, which now correspond to the FS-LKE. Thus, relation-specific separators (or prompts) have limited impact on factual knowledge estimation if subject-object pairs are correctly presented. Additionally, finding relation-specific prompts requires hand-crafted efforts or additional computation (Shin et al., 2020), unlike our zero-prompt many-shot approach using (subject, object) pairs. Therefore, ZP-LKE can potentially extend to any fact from knowledge graphs over any LLM, while HGPs and MMPs require additional supervision and relation-specific validation.
4.2. Evaluating Diverse Models and Relations
@@
We investigate the performance of 35 pre-trained LLMs and 14 fine-tuned LLMs across 50 relations using the ZP-LKE framework. Our analysis aims to uncover nuanced insights into the knowledge levels within these models. We will examine the results through two primary lenses: (1) the variations in knowledge across different model families, and (2) the influence of model size and fine-tuning within the same model family on their knowledge attributes.
4.2.1. Comparing different LLM families
@@
Some model families are consistently more knowledgeable than the rest. We sort the model families based on the performance of the model closest to 7B parameters 3337B parameters is a good reference point since all model families except GPT-NEO-X have models within a gap of 1B parameters: Mistral-7B, Gemma-7B, Llama-7B, Falcon-7B, MPT-7B, OPT-6.7B, GPT-J-6B, Pythia-6.9B, and Bloom-7.1B., and the models within each family based on average accuracy across 50 relations. Figure 7 shows that Mistral, Llama2, Gemma, and Llama families have higher performance on most of the relations than Pythia, Bloom, and OPT, indicating the latter’s lower factual knowledge.
@@
Different model families align in their relative factual knowledge. Although different model families have different knowledge levels, they have similar knowledge structures. We investigate the correlations between each model pair’s performance over 50 relations to assess the agreement in their knowledge levels. We compute the average correlations within each model family (e.g. Llama2 7B, 13B, 70B) in Figure 8. Despite differences in architecture and training datasets among model families, there is a significant consensus (correlation ¿ 0.6) regarding the hierarchy of knowledge across various relations. We also compile the three best and worst-performing relations for each model in Table 10, illustrating the consensus among all models. The consistent underperformance across specific relations also suggests that certain types of knowledge are universally less well-represented across different models, regardless of their architecture or size. This consistency in less-known knowledge across models highlights a potential vulnerability that could be exploited if these weaknesses are not addressed. Figure 16 shows the correlations between all the models within each family.
4.2.2. Comparing within the same LLM family